# Possible Liquid-Nitrogen-Temperature Superconductivity Driven by Perpendicular Electric Field in the Single-Bilayer Film of La3Ni2O7 at Ambient Pressure

### v2

@@ I.1 Consideration and a Simplified Study
We have further employed the DMRG approach, which can capture the quantum fluctuation beyond mean-field, to compute the ground state of Hamiltonian (1) under different electric fields $\varepsilon$ and the electron-doping levels of the $d_{x^{2}-y^{2}}$ orbitals $\delta=n_{tx}+n_{bx}-1$. For $\varepsilon=0$, we have $\delta=0$. When $\varepsilon$ increases, it drives electrons from $d_{z^{2}}$-orbitals in the top layer to $d_{x^{2}-y^{2}}$-orbitals in both layers, increasing $\delta$. However, as the exact relationship between $\varepsilon$ and $\delta$ is unclear, we set them as two independent variables in our DMRG study. The parameters $t_{\parallel}$ and $J_{\parallel}$ take the same values as the ones in the SBMF study while $\tilde{J}_{\perp}=0.8J_{\parallel}$ is adopted in the DMRG study. To characterize the pairing symmetry and strength, we analyze the interlayer pairing correlation function $\Phi^{\perp}(r)$ and the intra-bottom-layer pairing correlation function $\Phi^{\parallel}_{b}(r)$. More details are provided in METHODS.
Fig. 4 (a) shows the pairing phase diagram with respect to $\delta$ ($=0,1/16,1/8$) and $\varepsilon$ ($\in[0,1.6t_{\parallel}]$). Fig. 4 (b) shows the absolute value of the intra-bottom-layer pairing correlation functions $|\Phi^{\parallel}_{b}(r)|$ under different electric fields $\varepsilon=0,0.4t_{\parallel},0.8t_{\parallel},1.2t_{\parallel},1.6t_{\parallel}$ for $\delta=0$, and the results for $\delta=1/16$ are presented in Fig. 4 (c). It turns out that $|\Phi^{\parallel}_{b}(r)|$ exhibits algebraic decay under an non-zero external electric field with the decaying power exponent to be $K_{\text{SC}}$, i.e. $|\Phi^{\parallel}_{b}(r)|\propto r^{-K_{\text{SC}}}$ for large enough $r$, implying the presence of pairing within the bottom layers. Fig. 4 (d) and (e) depict $|\Phi^{\parallel}_{b}(r)|$ for different $d_{x^{2}-y^{2}}$-electron-doping levels $\delta=0,1/16,1/8$ under $\varepsilon=0.4t_{\parallel}$ in (d) and $\varepsilon=0.8t_{\parallel}$ in (e). All the algebraic decay exponents $K_{\text{SC}}$ are provided accordingly.
The results indicate that (i) With the enhancement of the perpendicular electric field $\varepsilon$, and hence the $d_{x^{2}-y^{2}}$-electron-doping level $\delta$, the pairing symmetry changes from interlayer $s$-wave to intra-bottom-layer $d$-wave (The criterion of the pairing symmetry is provided in METHODS); (ii) For all the $d_{x^{2}-y^{2}}$-electron-doping levels $\delta$ tested, the enhancement of the perpendicular electric field $\varepsilon$ leads to a reduction of $K_{\text{SC}}$, suggesting the enhancement of the intra-bottom-layer pairing; (iii) Under all the perpendicular electric field strengths $\varepsilon$ tested, the enhancement of the $d_{x^{2}-y^{2}}$-electron-doping level $\delta$ leads to a reduction of $K_{\text{SC}}$, suggesting the enhancement of the intra-bottom-layer pairing. From (ii) and (iii), it is clear that the enhancement of $\varepsilon$ and hence $\delta$ will significantly enhance the intra-bottom-layer pairing. These results are qualitatively consistent with those of the SBMF study. More results are given in the SI [146].
@@
Figure 4: The DMRG results. (a) The $\delta-\varepsilon$ phase diagram of the ground state. The red region corresponds to the $s$-wave pairing and the blue region to the $d$-wave pairing. (b)-(c) The absolute value of the intra-bottom-layer pairing correlation functions $|\Phi^{\parallel}_{b}(r)|$ under different electric fields $\varepsilon=0,0.4t_{\parallel},0.8t_{\parallel},1.2t_{\parallel},1.6t_{\parallel}$ for $\delta=0$ in (b) and $\delta=1/16$ in (c). (d)-(e) $|\Phi^{\parallel}_{b}(r)|$ for different $d_{x^{2}-y^{2}}$-electron-doping levels $\delta=0,1/16,1/8$ under $\varepsilon=0.4t_{\parallel}$ in (d) and $\varepsilon=0.8t_{\parallel}$ in (e). The algebraic decay exponents $K_{\text{SC}}$ are written in the four figures as well. In (a-e), $\delta$ and $\varepsilon$ are set as independent variables, since their exact relationship is unclear.
@@ I.2 The comprehensive two-orbital study
The above simplified single-orbital study has drawbacks: We cannot determine the relationship between the electron-doping of the $d_{x^{2}-y^{2}}$ orbitals and the electric field. In the SBMF study, we have to assume the ratio between the changes of the filling fractions of the two top-layer $E_{g}$ orbitals. In addition, we do not know how the neglected $3d_{z^{2}}$ orbital degree of freedom affects the pairing nature. To settle these puzzles, we conduct a comprehensive two-orbital model [107] to study with,
$H=$displaystyle-- $t_{\parallel}\sum_{\langle i,j\rangle,\mu}\left(c^{\dagger}_{i\mu x\sigma}c_{j\mu x\sigma}+\mathrm{h.c.}\right)-t_{\perp}\sum_{i}\left(c^{\dagger}_{itz\sigma}c_{ibz\sigma}+\mathrm{h.c.}\right)-t_{xz}\sum_{\langle i,j\rangle\mu}\left(c^{\dagger}_{i\mu x\sigma}c_{j\mu z\sigma}+(z\leftrightarrow x)+\mathrm{h.c.}\right)$ (2)
$+$ $J_{\parallel}\sum_{\langle i,j\rangle\mu}\bm{S}_{i\mu x}\cdot\bm{S}_{j\mu x}+J_{\perp}\sum_{i}\bm{S}_{itz}\cdot\bm{S}_{ibz}+\tilde{J}_{\perp}\sum_{i}\bm{S}_{itx}\cdot\bm{S}_{ibx}+\epsilon_{z}\sum_{i\mu\sigma}n_{i\mu z\sigma}+\epsilon_{x}\sum_{i\mu\sigma}n_{i\mu x\sigma}$
@@
The operators $c_{i\mu\alpha\sigma}$, $n_{i\mu\alpha}$, $\bm{S}_{i\mu\alpha}$ take the same meanings as those in model (1) except for an extra index $\alpha=x/z$ labeling the orbital. Note that $\bm{S}_{i\mu\alpha}$ for each orbital is spin-$\frac{1}{2}$ operator. $\epsilon_{\alpha}$ denotes the on-site energy of the orbital $\alpha$. We adpot the tight-binding (TB) parameters reported in Ref. [141], i.e. $t_{\parallel}=0.445\text{ eV}$, $t_{xz}=0.221\text{ eV}$, $t_{\perp}=0.503\text{ eV}$ and $\epsilon_{x}-\epsilon_{z}=0.367\text{ eV}$. The superexchange interactions are obtained through $J\approx 4t^{2}/U$, with $U=10t_{\parallel}$. Finally $\varepsilon$ denotes the voltage between the two layers. More details are provided in METHODS.
@@
The result shown in Fig. 5(b) for the comprehensive two-orbital study and that shown in Fig. 3(b) for the simplified one-orbital study look similar, except that in Fig. 5(b) the result is expressed as function of the directly controllable quantity $\varepsilon$. Actually, if we replace the $x$-axis of Fig. 5(b) by the calculated $n_{bx}=1-\delta_{bx}$, the resulting curve nearly coincides with Fig. 3(b), particularly in the large-$n_{bx}$ regime, see the SI [146]. The main reason for such similarity lies in that under strong $\bm{\varepsilon}$, the dominant superconducting pairing is the intra-bottom-layer $3d_{x^{2}-y^{2}}$-orbital pairing, which is insensitive to the $3d_{z^{2}}$ orbital. The main new information obtained in the two-orbital study lies in that the $3d_{z^{2}}$ orbital form interlayer $s$-wave pseudo-gap which is mixed with the intra-bottom-layer $d$-wave HTSC of the $3d_{x^{2}-y^{2}}$ orbital in the ratio of $1:\mathrm{i}$, as shown in Fig. 5(d). This state breaks time-reversal symmetry, although the experimentally detected superconducting gap is the standard $d$-wave gap of the $3d_{x^{2}-y^{2}}$ orbital. This intriguing result is left for experimental verification.
[…]
@@ III.1 The one-orbital model
The SBMF theory is used to solve the one-orbital model (1). In the SBMF approach, the superexchange terms are decomposed in $\chi-\Delta$ channel, e.g. $\bm{S}_{it}\cdot\bm{S}_{ib}=-\frac{3}{8}\left(\left\langle\chi^{\perp{\dagger}}\right\rangle\chi^{\perp}_{i}+\mathrm{h.c.}+\left\langle\Delta^{\perp{\dagger}}\right\rangle\Delta^{\perp}_{i}+\mathrm{h.c.}\right)$, $\chi$ and $\Delta$ represents hopping and pairing operators respectively. These MF parameters are further solved in a self-consistent manner. The specific steps can be referenced from prior work [79, 145, 107] and SI [146].
We also employ the DMRG method [147, 148] to solve the ground state of the Hamiltonian (1) as a comparison for the SBMF approach. The tensor libraries TensorKit [149] and FiniteMPS [150] provide an implementation of the required symmetry [151, 152]. We study the model on a $2\times 2\times L_{x}$ lattice with the open boundary conditions in the $x$ direction and choose $L_{x}=64$ for calculations. The matrix product state is constructed as shown in Fig. 6. We keep up to $D=12000$ $\mathrm{U}(1)_{\text{charge}}\times\mathrm{SU}(2)_{\text{spin}}$ multiplets in DMRG simulations and ensure the convergence accuracy of $10^{-6}$.
@@
$\Delta^{\parallel{\dagger}}_{i\mu}\equiv\Delta^{\bm{x}{\dagger}}_{i\mu}=\frac{1}{\sqrt{2}}\left(c^{{\dagger}}_{i\mu\uparrow}c^{{\dagger}}_{i+\bm{x},\mu\downarrow}-c^{{\dagger}}_{i\mu\downarrow}c^{{\dagger}}_{i+\bm{x},\mu\uparrow}\right),$
$\Delta^{\bm{y}{\dagger}}_{i\mu}=\frac{1}{\sqrt{2}}\left(c^{{\dagger}}_{i\mu\uparrow}c^{{\dagger}}_{i+\bm{y},\mu\downarrow}-c^{{\dagger}}_{i\mu\downarrow}c^{{\dagger}}_{i+\bm{y},\mu\uparrow}\right).$
Here, the subscripts $i+\bm{x}$($i+\bm{y}$) represent the NN site of the site $\bm{i}$ in the $x$($y$) direction. Fig. 7 shows how the singlet pairing operators are defined.
Figure 7: Illustration of the singlet pairing operators $\Delta^{\perp\dagger}_{i}$, $\Delta^{\parallel\dagger}_{i\mu}$ and $\Delta^{\bm{y}\dagger}_{i\mu}$.
@@
$\Phi^{\parallel}_{\mu}(r)\equiv\Phi^{\bm{xx}}_{\mu}(r)=\left\langle\Delta_{i\mu}^{\parallel{\dagger}}\Delta^{\parallel}_{j\mu}\right\rangle,$
$\Phi^{\bm{xy}}_{\mu}(r)=\left\langle\Delta_{i\mu}^{\parallel{\dagger}}\Delta^{\bm{y}}_{j\mu}\right\rangle,$
where $r=|\bm{i}-\bm{j}|$ is the distance between the sites $i$ and $j$.
@@
The dominant pairing channel is related to pairing symmetry. For the case where interlayer pairing dominates, the pairing symmetry is restricted to $s$-wave pairing; while for the case where intralayer pairing in the bottom-layer dominates, we determine the pairing symmetry by the sign function $\mathrm{sgn}\left[\Phi^{\parallel}_{b}(r)\Phi^{\bm{xy}}_{b}(r)\right]$. If $\mathrm{sgn}\left[\Phi^{\parallel}_{b}(r)\Phi^{\bm{xy}}_{b}(r)\right]=-1$ holds for all $r$, the ground state can be identified as the $d$-wave pairing SC state. See SI [146] for more details on DMRG.

### v3

@@ I.1 Consideration and a Simplified Study
We have further employed the DMRG approach, which can capture the quantum fluctuation beyond mean-field, to compute the ground state of Hamiltonian (1) under different electric fields $\varepsilon$ and the transferred electron-doping levels of the $d_{x^{2}-y^{2}}$ orbitals $\delta=n_{tx}+n_{bx}-1$. For $\varepsilon=0$, we have $\delta=0$. When $\varepsilon$ increases, it drives electrons from $d_{z^{2}}$-orbitals in the top layer to $d_{x^{2}-y^{2}}$-orbitals in both layers, increasing $\delta$. However, as the exact relationship between $\varepsilon$ and $\delta$ is unclear, we set them as two independent variables in our DMRG study. The parameters $t_{\parallel}$ and $J_{\parallel}$ take the same values as the ones in the SBMF study while $\tilde{J}_{\perp}=0.8J_{\parallel}$ is adopted in the DMRG study. To characterize the pairing symmetry and strength, we analyze the interlayer pairing correlation function $\Phi^{\perp}(r)$ and the intra-bottom-layer pairing correlation function $\Phi^{\parallel}_{b}(r)$. More details are provided in METHODS.
@@
Figure 4: The DMRG results. (a) The $\delta-\varepsilon$ phase diagram of the ground state. The red region corresponds to the $s$-wave pairing and the blue region to the $d$-wave pairing. (b)-(c) The absolute value of the intra-bottom-layer pairing correlation functions $|\Phi^{\parallel}_{b}(r)|$ under different electric fields $\varepsilon=0,0.4t_{\parallel},0.8t_{\parallel},1.2t_{\parallel},1.6t_{\parallel}$ for $\delta=0$ in (b) and $\delta=1/16$ in (c). (d)-(e) $|\Phi^{\parallel}_{b}(r)|$ for different transferred $d_{x^{2}-y^{2}}$-electron-doping levels $\delta=0,1/16,1/8$ under $\varepsilon=0.4t_{\parallel}$ in (d) and $\varepsilon=0.8t_{\parallel}$ in (e). The algebraic decay exponents $K_{\text{SC}}$ are written in the four figures as well, reflecting the decay rate of the pairing correlation function with spatial distance, negatively correlated with the corresponding pairing strength. In (a-e), $\delta$ and $\varepsilon$ are set as independent variables, since their exact relationship is unclear.
Fig. 4 (a) shows the pairing phase diagram with respect to $\delta$ (=0,1/16,1/8=0,1/16,1/8) and $\varepsilon$ ($\in[0,1.6t_{\parallel}]$). Fig. 4 (b) shows the absolute value of the intra-bottom-layer pairing correlation functions $|\Phi^{\parallel}_{b}(r)|$ under different electric fields $\varepsilon=0,0.4t_{\parallel},0.8t_{\parallel},1.2t_{\parallel},1.6t_{\parallel}$ for $\delta=0$, and the results for $\delta=1/16$ are presented in Fig. 4 (c). It turns out that $|\Phi^{\parallel}_{b}(r)|$ exhibits algebraic decay under an non-zero external electric field with the decaying power exponent to be $K_{\text{SC}}$, i.e. $|\Phi^{\parallel}_{b}(r)|\propto r^{-K_{\text{SC}}}$ for large enough $r$, implying the presence of pairing within the bottom layers. Fig. 4 (d) and (e) depict $|\Phi^{\parallel}_{b}(r)|$ for different transferred $d_{x^{2}-y^{2}}$-electron-doping levels $\delta=0,1/16,1/8$ under $\varepsilon=0.4t_{\parallel}$ in (d) and $\varepsilon=0.8t_{\parallel}$ in (e). All the algebraic decay exponents $K_{\text{SC}}$ are provided accordingly.
The results indicate that (i) With the enhancement of the perpendicular electric field $\varepsilon$, and hence the transferred $d_{x^{2}-y^{2}}$-electron-doping level $\delta$, the pairing symmetry changes from interlayer $s$-wave to intra-bottom-layer $d$-wave (The criterion of the pairing symmetry is provided in METHODS); (ii) For all the transferred $d_{x^{2}-y^{2}}$-electron-doping levels $\delta$ tested, the enhancement of the perpendicular electric field $\varepsilon$ leads to a reduction of $K_{\text{SC}}$, suggesting the enhancement of the intra-bottom-layer pairing; (iii) Under all the perpendicular electric field strengths $\varepsilon$ tested, the enhancement of the transferred $d_{x^{2}-y^{2}}$-electron-doping level $\delta$ leads to a reduction of $K_{\text{SC}}$, suggesting the enhancement of the intra-bottom-layer pairing. From (ii) and (iii), it is clear that the enhancement of $\varepsilon$ and hence $\delta$ will significantly enhance the intra-bottom-layer pairing. These results are qualitatively consistent with those of the SBMF study. More results are given in the SI [SM].
Besides, we study the effect of interlayer Coulomb interaction. Our results show that the interlayer Coulomb interaction slightly promotes charge transfer between layers and the intra-bottom-layer pairing, while suppressing the interlayer pairing. See SI [SM] for details.
@@ I.2 The comprehensive two-orbital study
The above simplified single-orbital study has drawbacks: We cannot determine the relationship between the electron-doping of the $d_{x^{2}-y^{2}}$ orbitals and the electric field. In the SBMF study, we have to assume the ratio between the changes of the filling fractions of the two top-layer $E_{g}$ orbitals. In addition, we do not know how the neglected $3d_{z^{2}}$ orbital degree of freedom affects the pairing nature. To settle these puzzles, we conduct a comprehensive two-orbital model [Lu2024interplay] to study with,
$H=$ - $t_{\parallel}\sum_{\langle i,j\rangle,\mu}\hat{P}\left(c^{\dagger}_{i\mu x\sigma}c_{j\mu x\sigma}+\mathrm{h.c.}\right)\hat{P}-t_{\perp}\sum_{i}\hat{P}\left(c^{\dagger}_{itz\sigma}c_{ibz\sigma}+\mathrm{h.c.}\right)\hat{P}-t_{xz}\sum_{\langle i,j\rangle\mu}\hat{P}\left(c^{\dagger}_{i\mu x\sigma}c_{j\mu z\sigma}+(z\leftrightarrow x)+\mathrm{h.c.}\right)\hat{P}$ (2)
$+$ $J_{\parallel}\sum_{\langle i,j\rangle\mu}\mathbf{S}_{i\mu x}\cdot\mathbf{S}_{j\mu x}+J_{\perp}\sum_{i}\mathbf{S}_{itz}\cdot\mathbf{S}_{ibz}+\tilde{J}_{\perp}\sum_{i}\mathbf{S}_{itx}\cdot\mathbf{S}_{ibx}+\epsilon_{z}\sum_{i\mu\sigma}n_{i\mu z\sigma}+\epsilon_{x}\sum_{i\mu\sigma}n_{i\mu x\sigma}$
@@
The operators $c_{i\mu\alpha\sigma}$, $n_{i\mu\alpha}$, $\mathbf{S}_{i\mu\alpha}$ take the same meanings as those in model (1) except for an extra index $\alpha=x/z$ labeling the orbital, and $\hat{P}$ is a projection operator projecting out the double occupancy in the same orbital of all sites. Note that $\mathbf{S}_{i\mu\alpha}$ for each orbital is spin-$\frac{1}{2}$ operator. $\epsilon_{\alpha}$ denotes the on-site energy of the orbital $\alpha$. We adpot the tight-binding (TB) parameters reported in Ref. [Daoxin_Yao2025], i.e. $t_{\parallel}=0.445\text{ eV}$, $t_{xz}=0.221\text{ eV}$, $t_{\perp}=0.503\text{ eV}$ and $\epsilon_{x}-\epsilon_{z}=0.367\text{ eV}$. The superexchange interactions are obtained through $J_{\parallel}\approx 4t_{\parallel}^{2}/U$ and $J_{\perp}\approx 4t_{\perp}^{2}/U$, with $U=10t_{\parallel}$. Finally $\varepsilon$ denotes the voltage between the two layers. Here, due to the weak super-exchange interaction between the $d_{z^{2}}$ orbitals in the layer, we do not consider this term in our model. In addition, the Hund’s coupling $J_{H}$ of La3Ni2O7 is generally considered to be in the range of 0.7 eV to 1 eV in past studies [lechermann2023, ouyang2023hund, WuWei2023charge], which only slightly larger than the largest hopping parameter $t_{\perp}=0.503~\mathrm{eV}$ and thus does not satisfy the premise of the Schriffer-Wolf transformation or perturbation theory, we do not apply it here. More details are provided in METHODS.
Our SBMF results of Eq. (2) (see METHODS and the SI [SM]) are shown in Fig. 5. Fig. 5(a) shows the $\varepsilon$-dependence of the hole densities $\delta_{\mu\alpha}$. Obviously, the $\delta_{tz}$ enhances obviously with $\varepsilon$, suggesting that the top-$3d_{z^{2}}$ orbital is donating electrons. These donated electrons flow to the $3d_{x^{2}-y^{2}}$ orbitals in both layers, with more of them flowing to the bottom layer when $\varepsilon>0.1\text{ eV}$ while about half of them flow to the bottom layer when $\varepsilon\leq 0.1\text{ eV}$. Fig. 5(b) shows the $\varepsilon$-dependence of the pairing symmetry and the pairing gap amplitude of the bottom-layer $3d_{x^{2}-y^{2}}$ orbital. At low $\varepsilon\leq 0.03$ eV, the pairing symmetry is $s$-wave, whose pairing configuration is shown in Fig. 5(c), wherein the $3d_{z^{2}}$-orbital form interlayer $s$-wave pseudo-gap, while the $3d_{x^{2}-y^{2}}$ orbital form $s$-wave SC with coexisting intralayer and interlayer pairing. In this regime the interlayer pairing is suppressed by the enhancement of $\varepsilon$ while the intralayer pairing is enhanced. When $\varepsilon=0$, the interlayer pairing gap is the largest. When $\varepsilon$ is about $0.01\sim 0.03\text{ eV}$, the intralayer pairing gap is slightly larger than the interlayer pairing gap. When $\varepsilon>0.03$ eV, the pairing symmetry is dd($3d_{x^{2}-y^{2}}$)+iss($3d_{z^{2}}$), whose pairing configuration is shown in Fig. 5(d). In this state, the $3d_{z^{2}}$ orbital form interlayer $s$-wave pseudo-gap, while the bottom-layer $3d_{x^{2}-y^{2}}$ orbital form intralayer $d$-wave SC. When $\varepsilon$ enhances in this regime, the pairing amplitude for the $d$-wave part enhances promptly. For $\varepsilon>0.13$ eV, the pairing amplitude can go beyond $0.02$ eV. Then from the relation $T_{c}\approx 0.42\tilde{\Delta}$ for the $d$-wave SC illustrated in Fig. 3(b), we have got HTSC with $T_{c}\gtrsim 80$ K!
@@
Figure 5: The SBMF results for the two-orbital model. (a) The hole densities $\delta_{\mu\alpha}$ for the three orbitals as functions of the strength of the electric field $\varepsilon$. (b) The pairing gap amplitude of the bottom-layer $3d_{x^{2}-y^{2}}$-orbital as function of $\varepsilon$. (c)-(d) The pairing configurations of the $s$-wave and the dd($3d_{x^{2}-y^{2}}$)+iss($3d_{z^{2}}$)-wave, respectively.
@@ III.1 The one-orbital model
The SBMF theory is used to solve the one-orbital model (1). In the SBMF approach, the superexchange terms are decomposed in $\chi-\Delta$ channel, e.g. $\mathbf{S}_{it}\cdot\mathbf{S}_{ib}=-\frac{3}{8}\left(\left\langle\chi^{\perp{\dagger}}\right\rangle\chi^{\perp}_{i}+\mathrm{h.c.}+\left\langle\Delta^{\perp{\dagger}}\right\rangle\Delta^{\perp}_{i}+\mathrm{h.c.}\right)$, $\chi$ and $\Delta$ represents hopping and pairing operators respectively. These MF parameters are further solved in a self-consistent manner. The specific steps can be referenced from prior work [lu2023bilayertJ, kotliar1988, Lu2024interplay] and SI [SM].
We also employ the DMRG method [white1993dmrg, weng1999dmrg] to solve the ground state of the Hamiltonian (1) as a comparison for the SBMF approach. The tensor libraries TensorKit [jutho_2024_13950435] and FiniteMPS [Li_FiniteMPS_jl_2024] provide an implementation of the required symmetry [WEICHSELBAUM20122972, PhysRevResearch.2.023385]. We study the model on a $2\times 2\times L_{x}$ lattice with the open boundary conditions in the $x$ direction and choose $L_{x}=64$ for calculations. The matrix product state is constructed as shown in Fig. 6. We keep up to $D=12000$ $\mathrm{U}(1)_{\text{charge}}\times\mathrm{SU}(2)_{\text{spin}}$ multiplets in DMRG simulations and ensure the convergence accuracy of $10^{-6}$.
@@
$\Delta^{\parallel{\dagger}}_{i\mu}\equiv\Delta^{\mathbf{x}{\dagger}}_{i\mu}=\frac{1}{\sqrt{2}}\left(c^{{\dagger}}_{i\mu\uparrow}c^{{\dagger}}_{i+\mathbf{x},\mu\downarrow}-c^{{\dagger}}_{i\mu\downarrow}c^{{\dagger}}_{i+\mathbf{x},\mu\uparrow}\right),$
$\Delta^{\mathbf{y}{\dagger}}_{i\mu}=\frac{1}{\sqrt{2}}\left(c^{{\dagger}}_{i\mu\uparrow}c^{{\dagger}}_{i+\mathbf{y},\mu\downarrow}-c^{{\dagger}}_{i\mu\downarrow}c^{{\dagger}}_{i+\mathbf{y},\mu\uparrow}\right).$
Here, the subscripts $i+\mathbf{x}$($i+\mathbf{y}$) represent the NN site of the site $\mathbf{i}$ in the xx($y$) direction. Fig. 7 shows how the singlet pairing operators are defined.
Figure 7: Illustration of the singlet pairing operators $\Delta^{\perp\dagger}_{i}$, $\Delta^{\parallel\dagger}_{i\mu}$ and $\Delta^{\mathbf{y}\dagger}_{i\mu}$.
@@
$\Phi^{\parallel}_{\mu}(r)\equiv\Phi^{\mathbf{xx}}_{\mu}(r)=\left\langle\Delta_{i\mu}^{\parallel{\dagger}}\Delta^{\parallel}_{j\mu}\right\rangle,$
$\Phi^{\mathbf{xy}}_{\mu}(r)=\left\langle\Delta_{i\mu}^{\parallel{\dagger}}\Delta^{\mathbf{y}}_{j\mu}\right\rangle,$
where $r=|\mathbf{i}-\mathbf{j}|$ is the distance between the sites $i$ and $j$.
@@
The dominant pairing channel is related to pairing symmetry. For the case where interlayer pairing dominates, the pairing symmetry is restricted to $s$-wave pairing; while for the case where intralayer pairing in the bottom-layer dominates, we determine the pairing symmetry by the sign function $\mathrm{sgn}\left[\Phi^{\parallel}_{b}(r)\Phi^{\mathbf{xy}}_{b}(r)\right]$. If $\mathrm{sgn}\left[\Phi^{\parallel}_{b}(r)\Phi^{\mathbf{xy}}_{b}(r)\right]=-1$ holds for all $r$, the ground state can be identified as the $d$-wave pairing SC state. See SI [SM] for more details on DMRG.